# Attainability of quantum state discrimination bounds with collective measurements on finite copies

### v1

@@ III.1.3 Chernoff bound
which allows the quantity $\epsilon_{\infty}$ to be calculated.
III.1.4 Comparison of bounds
We are now in a position to evaluate the quantities $R(M)$ and $\delta(M,N)$. In Fig. 2 a), we plot $1-R(M)$ as a function of $M$. Although it is not possible to say definitively that $R(M)$ does not plateau to 1 for any finite $M$, it certainly appears that way from this figure. In Fig. 2 b), we plot $\delta(M+2,M)$ as a function of $M$. In this case it is evident that this function is continuously decreasing. However, again an analytic proof of this statement is not presented here.
As this example is essentially classical (the two states are simultaneously diagonalisable), this appears to suggest that the inability to exactly saturate the Chernoff bound is not a quantum property. Indeed similar concepts have been studied in classical information theory, where second order corrections to the asymptotic limits are known hayashi2009information .
For reference, we also present $R_{\epsilon}(M)$ and $\delta_{\epsilon}(M,M+2)$ in appendix C.
III.1.5 Exponent attained by repeatedly performing the Helstrom measurement
@@
While the above comparison is useful for examining how the Helstrom bound tends to the Chernoff bound, this comparison is somewhat irrelevant for small numbers of copies of the unknown state. This is because the Chernoff bound is most relevant in the asymptotic limit. As such in this section, we consider what happens when we have $N$ total copies of the quantum state and we perform the $M$-copy Helstrom measurement $N/M$ times, assuming $N\gg M$. A decision on which state is present can then be made by majority vote. Physically this corresponds to the situation where we receive either the state $\rho_{+}^{\otimes N}$ or $\rho_{-}^{\otimes N}$, however we are only capable of performing an entangling measurement on $M\ll N$ copies. It should be noted that repeatedly performing the Helstrom measurement is not the optimal strategy when allowing for $M$-copy entangling measurements on $N$ total copies of the state. As such, we do not necessarily expect the error exponent from this strategy to tend to the Chernoff exponent in the asymptotic limit.
@@
The details of how the error exponent is calculated for this measurement strategy are presented in appendix D. In brief, the outcome of each Helstrom measurement is treated as a binomial random variable. We then make a final decision depending on which outcome, $\rho_{+}^{\otimes M}$ or $\rho_{-}^{\otimes M}$, occurs more frequently. This allows us to place bounds on the error exponent using the classical Chernoff bound. This is shown in Fig. 3. Interestingly we see that the single-copy Helstrom measurement appears to be sufficient to saturate the Chernoff bound in the asymptotic limit. In Fig. 3 b), we show the convergence of this strategy to the Chernoff bound. This is consistent with other known results, see for example Ref. acin2005multiple , where repeated single-copy measurements and a unanimity vote strategy was shown to saturate the Chernoff bound asymptotically. It is an interesting question, however not one that we address here, as to the optimal $M$-copy measurement for minimising the asymptotic error exponent. It may be that, for example, an optimised two-copy measurement can approach the Chernoff bound faster.
@@ III.2.2 Pure state Helstrom bound
First note that the Chernoff bound for pure states is given by $\text{cos}(\alpha)^{2M}$. For pure states, $\rho_{+}$ and $\rho_{-}$ only have one eigenvector and eigenvalue. Writing $\rho_{\pm}(v=0)=\ket{\psi_{\pm}}\bra{\psi_{\pm}}$ |, these are given by
@@
We can therefore evaluate the normalised ratio of the Helstrom bound to the Chernoff bound (Eq. (12)) as
In appendix E we prove that this ratio never reaches the value 1 for any finite $M$, hence the Chernoff bound cannot be exactly saturated. This is consistent with existing results for asymmetric quantum state discrimination pereira2023analytical ; audenaert2012quantum ; rouze2017finite ; li2014second . Of course it is still possible that the Chernoff bound can be approximately saturated, as is shown in Fig. 4. In this figure, we plot 1 minus the normalised ratio of the $M$-copy error probability to the asymptotic error probability. This plot shows that the ratio is an increasing function of $M$, however the ratio never reaches exactly 1. Therefore, the asymptotic limit cannot be reached with any finite-copy measurement. This result is stated as the following theorem
Theorem 1
III.2.3 Numerical simulations for mixed states
Figure 6: Scaling of the error probability for a quantum metrology inspired example. a) We plot the difference between the exact Helstrom bound and the approximation based on Eq. (34) as a function of $\theta$. b) We show 1 minus the normalised ratio of the $M$-copy error probability using the approximation in Eq. (34), to the asymptotic error exponent.
III.3 Example 3 - A quantum metrology inspired example
Let us now consider a third qubit problem - distinguishing between $\rho_{1}=\mathbb{I}/2$ and $\rho_{2}=\mathbb{I}/2+\theta_{x}\sigma_{x}+\theta_{y}\sigma_{y}+\theta_{z}\sigma_{z}$ where $\theta_{x},\theta_{y},\theta_{z}\ll 1$. This was considered in Ref. li2023optimal , in a quantum metrology setting of trying to estimate the parameters $\theta$. From the gap persistence theorem in quantum metrology conlon2022gap , we know that for estimating the parameters $\theta$, the asymptotic limits on the attainable precision cannot be reached with a measurement on any finite number of copies. Thus, one might intuitively expect a similar result for distinguishing $\rho_{1}$ and $\rho_{2}$.
III.3.1 Helstrom bound
In the single-copy case
$\rho_{2}-\rho_{1}=\theta_{x}\sigma_{x}+\theta_{y}\sigma_{y}+\theta_{z}\sigma_{z}$ (30)
Therefore
$\begin{split}P^{\text{H}}_{\text{e}}(1)&=\frac{1}{2}\bigg{(}1-\left\lVert\theta_{x}\sigma_{x}+\theta_{y}\sigma_{y}+\theta_{z}\sigma_{z}\right\rVert\bigg{)}\\ &=\frac{1}{2}\bigg{(}1-\sqrt{\theta_{x}^{2}+\theta_{y}^{2}+\theta_{z}^{2}}\bigg{)}\;.\end{split}$ (31)
Now let us consider the $M$-copy situation. Dropping terms that are second order or higher in $\theta$ and writing $\theta A=\theta_{x}\sigma_{x}+\theta_{y}\sigma_{y}+\theta_{z}\sigma_{z}$, we see
$\rho_{2}^{\otimes M}\approx\frac{\mathbb{I}_{2^{M}}}{2^{M}}+\frac{1}{2^{M-1}}\sum_{i}\theta A^{(i)}\;,$ (32)
where we use $\mathbb{I}_{x}$ to denote the identity matrix in an $x$-dimensional Hilbert space and $A^{(i)}=\mathbb{I}\otimes\ldots\mathbb{I}\otimes A\otimes\mathbb{I}\ldots\mathbb{I}$ with $A$ in the $i$th position. Similar to the example in section III.1, we can then rewrite the matrix $\theta A$ in a different basis as
$\theta A=\begin{pmatrix}\sqrt{\theta_{x}^{2}+\theta_{y}^{2}+\theta_{z}^{2}}&0\\ 0&-\sqrt{\theta_{x}^{2}+\theta_{y}^{2}+\theta_{z}^{2}}\end{pmatrix}\;.$ (33)
Denoting $a=\sqrt{\theta_{x}^{2}+\theta_{y}^{2}+\theta_{z}^{2}}$, we can then see that
$\norm{\rho_{2}^{\otimes M}-\rho_{1}^{\otimes M}}\approx\frac{a}{2^{M-2}}\sum_{i=0}^{\lceil M/2\rceil}\binom{M}{i-1}(M-2(i-1))\;.$ (34)
This allows an approximate expression for the Helstrom bound to be obtained. In Fig. 6 a) we plot the difference between this approximation and the true Helstrom bound, which we label as $\Delta(\theta)$. In this simulation we assume $\theta=\theta_{x}=\theta_{y}=\theta_{z}$. As expected the approximation gets worse as both $\theta$ and $M$ increase.
III.3.2 Chernoff bound
We can rewrite $\rho_{2}$ in a different basis as
$\rho_{2}=\mathbb{I}/2+\begin{pmatrix}a&0\\ 0&-a\end{pmatrix}\;.$ (35)
For example when $\theta_{x}=\theta_{y}=0$, $a=\theta_{z}$ Then the Chernoff bound is optimised when
$s=\frac{-\mathrm{i}\pi+\text{log}(\text{l}_{-})-\text{log}(\text{l}_{+})+\text{log}(\text{log}(\text{l}_{-}))-\text{log}(\text{log}(\text{l}_{+}))}{\text{log}(\text{l}_{-})-\text{log}(\text{l}_{+})}\;.$ (36)
where $\text{l}_{\pm}=1\pm 2a$. Note that the $-\mathrm{i}\pi$ term ensures that the solution is real overall. Again this lets us make a comparison of the two bounds. We do this in Fig. 6 b). In spite of the error in our approximate solution to the Helstrom bound (Fig. 6 a)), we still observe the expected convergence of the Helstrom error probability to the asymptotic error probability.
@@ IV Requirements to saturate the $M$-copy Helstrom bound
We now wish to examine the requirements for saturating the $M$-copy Helstrom bound for any finite value of $M$. In Table 2 we have summarised known results on when LOCC measurements can saturate the $M$-copy Helstrom bound. However, there exist other measurements which are more powerful than LOCC measurements but less powerful than a collective measurement on all $M$ copies. For the example considered in section III.2, we have previously shown that, with equal priors ($q=1/2$), the two-copy Helstrom bound can be saturated with LOCC measurements conlon2023discriminating , i.e. 2-copy collective measurements are not required. From a practical viewpoint, it is interesting to check whether this trend holds in general, i.e. are $M-1$-copy collective measurements sufficient to saturate the $M$-copy Helstrom bound? Or are there scenarios where, for example, different collective measurements on $M/2$ copies of the state can saturate the $M$-copy Helstrom bound?
Throughout this section, we will consider a class of measurements on $M$ copies of the quantum state, where the measurement is broken into 2 stages. In the first stage we perform a collective measurement on $M_{1}$ copies of the quantum state and in the second stage we perform a collective measurement, which depends on the results of the first measurement, on the remaining $M_{2}=M-M_{1}$ copies of the quantum state. We denote the error probability for this strategy as $P_{M_{1},M_{2}}^{M}$.
@@
and $\bra{\psi}\ket{\psi^{\perp}}=0$ higgins2011multiple . Denote the outcome of this measurement as $D_{i}$, so that $D_{i}$ can correspond to either $\Pi_{0}$ or $\Pi_{1}$ clicking. Given the measurement outcome $D_{i}$, the posterior probability after performing this single qubit measurement is given by
@@
Let us first return to the example considered in section III.1. As the two states in this example are simultaneously diagonalisable, the Helstrom bound can be saturated through the maximum likelihood measurement, see Ref. audenaert2014upper and appendix A.
@@
$\begin{split}\text{Tr}[\rho\Pi_{0}]&=\frac{1}{2}\big{(}1+(1-v)\text{cos}(\alpha)\text{cos}(2\theta)\big{)}\\ \text{Tr}[\rho\Pi_{1}]&=\frac{1}{2}\big{(}1-(1-v)\text{cos}(\alpha)\text{cos}(2\theta)\big{)}\;.\end{split}$ (40)
@@
$\begin{split}\text{Tr}[\rho_{+}\Pi_{0}]&=\frac{1}{4}\big{(}2+(1-v)\text{cos}(\alpha-2\theta)\big{)}\\ \text{Tr}[\rho_{+}\Pi_{1}]&=\frac{1}{4}\big{(}2-(1-v)\text{cos}(\alpha-2\theta)\big{)}\;.\end{split}$ (41)
@@
Interestingly we see that for $M=2$ and equal prior probabilities of the two states ($q=1/2$), LOCC measurements are sufficient to saturate the two-copy Helstrom bound. In appendix F, we prove this for all $\alpha$ and $v$. However, for $M=3$ and $M=4$, two- and three-copy collective measurements respectively, do not appear to be sufficient. This is shown in Fig. 7. Another strategy which we can analytically obtain results for, is to perform the $M-1$-copy Helstrom measurement on the first $M-1$ copies and then perform an optimised single qubit measurement on the final copy. Note however, that in this case the $M-1$-copy Helstrom measurement may not be the optimal measurement to minimise the overall error probability. Also note that the optimal measurement on the final copy will simply be the Helstrom measurement for the updated posterior probability. We denote the error probability for this strategy as $P_{H(M-1),1}^{M}$. For $M=3$ and $M=4$ these results are also shown in Fig. 7. Interestingly, even the LOCC measurements can outperform this particular measurement. Note that as there are two possible outcomes for the Helstrom measurement there are two single qubit measurement angles to be optimised in the final measurement stage. However, due to the symmetry of the problem, and for clarity of presentation, we simply plot twice the error probability for one of these outcomes as a function of the corresponding measurement angle. It should also be noted that for the $M=4$ case shown in Fig. 7, the optimal strategy is simple to make a guess based on the measurement of the first 3 copies. This is reflected in the fact that the $P_{H(3),1}^{4}$ strategy is not a function of the single qubit measurement angle.
@@
Figure 8: Attainability of the $M$-copy Helstrom bound with different measurement strategies. On the left we plot the probability of error for different measurement strategies optimised using parameterised quantum circuits. We see that no measurement strategy which does not use collective measurements on all 4-copies simultaneously can saturate the Helstrom bound. On the right we extend these results beyond $M=4$. $v=0.1$ and $\alpha=\pi/4$ for this figure.
IV.3 Numerical search for measurement strategies saturating the Helstrom bound
@@
As is noted in Table 2, for pure states ($v=0$) the Helstrom bound can be saturated by LOCC measurements brody1996minimum ; ban1997accessible ; acin2005multiple . We verify this numerically for our problem, confirming that the LOCC strategy can saturate the 4-copy Helstrom bound up to 5 significant figures.

### v2

@@ III.1.3 Chernoff bound
which allows the quantity $\epsilon_{\infty}$ (i.e. the asymptotic error exponent) to be calculated.
III.1.4 Exponent attained by repeatedly performing the Helstrom measurement
@@
In this section, we consider the scenario where we have $N$ total copies of the quantum state and we perform the $M$-copy Helstrom measurement $N/M$ times, assuming $N\gg M$. A decision on which state is present can then be made by majority vote. Physically this corresponds to the situation where we receive either the state $\rho_{+}^{\otimes N}$ or $\rho_{-}^{\otimes N}$, however we are only capable of performing an entangling measurement on $M\ll N$ copies. It should be noted that repeatedly performing the Helstrom measurement is not necessarily the optimal strategy when allowing for $M$-copy entangling measurements on $N$ total copies of the state. As such, we do not necessarily expect the error exponent from this strategy to tend to the Chernoff exponent in the asymptotic limit.
@@
The details of how the error exponent is calculated for this measurement strategy are presented in appendix C. In brief, the outcome of each Helstrom measurement is treated as a binomial random variable. We then make a final decision depending on which outcome, $\rho_{+}^{\otimes M}$ or $\rho_{-}^{\otimes M}$, occurs more frequently. This allows us to place bounds on the error exponent using the classical Chernoff bound. This is shown in Fig. 1. Interestingly we see that the single-copy Helstrom measurement appears to be sufficient to saturate the Chernoff bound in the asymptotic limit. In Fig. 1 b), we show the convergence of this strategy to the Chernoff bound. This is consistent with other known results, see for example Ref. [32], where repeated single-copy measurements and a unanimity vote strategy was shown to saturate the Chernoff bound asymptotically. It is an interesting question, however not one that we address here, as to the optimal $M$-copy measurement for minimising the asymptotic error exponent. It may be that, for example, an optimised two-copy measurement can approach the Chernoff bound faster.
Figure 2: Error exponent attained by repeatedly implementing the Helstrom measurement for the example in section III.2. a) The black line represents the error exponent given by the Chernoff bound. The green (odd $M$) and red (even $M$) points represent the true error probability obtained when repeatedly implementing the $M$-copy Helstrom bound for $N=70,000$. The grey lines and black circles correspond to the lower bound on the error exponent given by Eq. (33). b) The error exponents obtained when repeatedly implementing the single-copy Helstrom measurement are shown in green. The Chernoff bound is shown in black. Both plots use $v=0.8$ and $\alpha=\pi/4$.
@@ III.2.2 Pure state Helstrom bound
For pure states, $\rho_{+}$ and $\rho_{-}$ only have one eigenvector and eigenvalue. Writing $\rho_{\pm}(v=0)=\ket{\psi_{\pm}}\bra{\psi_{\pm}}$ |, these are given by
@@
III.2.3 Error exponent obtained by repeatedly performing the Helstrom measurement
For mixed states, we are able to compute the Helstrom bound numerically. This enables us to use the same approach as in section III.1.4, to evaluate how the Chernoff bound is approached as we repeatedly implement the Helstrom measurement. As is shown in Fig. 2 a), in this case the lower bound from appendix C is not as close to the true error exponent as in Fig. 1. However, as is shown in in Fig. 2 b), once again we see that repeatedly implementing the single-copy Helstrom bound is sufficient to asymptotically approach the Chernoff bound.
@@ IV Requirements to saturate the $M$-copy Helstrom bound
We now wish to examine the requirements for saturating the $M$-copy Helstrom bound for any finite value of $M$. In Table 2 we have summarised known results on when LOCC measurements can saturate the $M$-copy Helstrom bound. However, there exist other measurements which are more powerful than LOCC measurements but less powerful than a collective measurement on all $M$ copies. For the example considered in section III.2, we have previously shown that, with equal priors ($q=1/2$), the two-copy Helstrom bound can be saturated with LOCC measurements [14], i.e. 2-copy collective measurements are not required. From a practical viewpoint, it is interesting to check whether this trend holds in general, i.e. are $M-1$-copy collective measurements sufficient to saturate the $M$-copy Helstrom bound? Or are there scenarios where, for example, different collective measurements on $M/2$ copies of the state can saturate the $M$-copy Helstrom bound?
Throughout this section, we will consider a class of measurements on $M$ copies of the quantum state, where the measurement is broken into 2 stages. In the first stage we perform a collective measurement on $M_{1}$ copies of the quantum state and in the second stage we perform a collective measurement, which depends on the results of the first measurement, on the remaining $M_{2}=M-M_{1}$ copies of the quantum state. We denote the error probability for this strategy as $P_{M_{1},M_{2}}^{M}$. By introducing this family of measurements we are effectively constructing a bipartite quantum state discrimination problem, where the first party has $M_{1}$ copies of the quantum state and the second party has $M_{2}$ copies. We allow one-way LOCC measurements for the two parties. In this setting it is known that for certain examples LOCC measurements are insufficient to saturate the Helstrom bound - a phenomena known as non-locality without entanglement [54, 55].
@@
and $\bra{\psi}\ket{\psi^{\perp}}=0$ [39]. Denote the outcome of this measurement as $D_{i}$, so that $D_{i}$ can correspond to either $\Pi_{0}$ or $\Pi_{1}$ clicking. Given the measurement outcome $D_{i}$, the posterior probability after performing this single qubit measurement is given by
@@
Let us first return to the example considered in section III.1. As the two states in this example are simultaneously diagonalisable, the Helstrom bound can be saturated through the maximum likelihood measurement, see Ref. [36] and appendix A.
@@
$\begin{split}\text{Tr}[\rho\Pi_{0}]&=\frac{1}{2}\big{(}1+(1-v)\text{cos}(\alpha)\text{cos}(2\phi)\big{)}\\ \text{Tr}[\rho\Pi_{1}]&=\frac{1}{2}\big{(}1-(1-v)\text{cos}(\alpha)\text{cos}(2\phi)\big{)}\;.\end{split}$ (22)
@@
$\begin{split}\text{Tr}[\rho_{+}\Pi_{0}]&=\frac{1}{2}\big{(}1+(1-v)\text{cos}(\alpha-2\phi)\big{)}\\ \text{Tr}[\rho_{+}\Pi_{1}]&=\frac{1}{2}\big{(}1-(1-v)\text{cos}(\alpha-2\phi)\big{)}\;.\end{split}$ (23)
@@
Interestingly we see that for $M=2$ and equal prior probabilities of the two states ($q=1/2$), LOCC measurements are sufficient to saturate the two-copy Helstrom bound. In appendix D, we prove this for all $\alpha$ and $v$. However, for $M=3$ and $M=4$, two- and three-copy collective measurements respectively, do not appear to be sufficient. This is shown in Fig. 3. Another strategy which we can analytically obtain results for, is to perform the $M-1$-copy Helstrom measurement on the first $M-1$ copies and then perform an optimised single qubit measurement on the final copy. Note however, that in this case the $M-1$-copy Helstrom measurement may not be the optimal measurement to minimise the overall error probability. Also note that the optimal measurement on the final copy will simply be the Helstrom measurement for the updated posterior probability. We denote the error probability for this strategy as $P_{H(M-1),1}^{M}$. For $M=3$ and $M=4$ these results are also shown in Fig. 3. Interestingly, even the LOCC measurements can outperform this particular measurement. Note that as there are two possible outcomes for the Helstrom measurement there are two single qubit measurement angles to be optimised in the final measurement stage. However, due to the symmetry of the problem, and for clarity of presentation, we simply plot twice the error probability for one of these outcomes as a function of the corresponding measurement angle. It should also be noted that for the $M=4$ case shown in Fig. 3, the optimal strategy is simple to make a guess based on the measurement of the first 3 copies. This is reflected in the fact that the $P_{H(3),1}^{4}$ strategy is not a function of the single qubit measurement angle.
@@
IV.3 Variational quantum circuits for finding optimal bipartite LOCC measurement strategies
@@
Figure 4: Attainability of the $M$-copy Helstrom bound with different LOCC measurement strategies. On the left we plot the probability of error for different measurement strategies optimised using parameterised quantum circuits. We see that no measurement strategy which does not use collective measurements on all 4-copies simultaneously can saturate the Helstrom bound. On the right we extend these results beyond $M=4$. $v=0.1$ and $\alpha=\pi/4$ for this figure.
@@
As is noted in Table 2, for pure states ($v=0$) the Helstrom bound can be saturated by LOCC measurements [30, 31, 32]. We verify this numerically for our problem, confirming that the LOCC strategy can saturate the 4-copy Helstrom bound up to 5 significant figures.
@@
IV.4 Numerical search for other measurements saturating the Helstrom bound
Figure 5: Attainability of the $M$-copy Helstrom bound with different PPT measurement strategies. As in Fig. 4, however now we consider PPT measurements between a bipartition $M_{1},M_{2}$ as opposed to one-way LOCC measurements.
Figure 6: Attainability of the Helstrom bound with different PPT measurement strategies. In both figures the dashed line corresponds to a $y=x$ line. The left figure shows 10,000 randomly sampled problems discriminating between four copies of three known qubit states. The red and blue data points correspond to $P_{2,2}^{4,\text{ PPT}}$ and $P_{1,3}^{4,\text{ PPT}}$ measurements respectively. The inset shows the difference between the Helstrom measurement and the corresponding PPT measurement. The x-axis of the inset is the same as the main figure. The right figure shows 10,000 randomly sampled problems discriminating between three quantum states in a 16 dimensional Hilbert space.
Above we effectively considered a bipartite state discrimination problem with one-way LOCC measurements allowed between the two parties. To further examine the attainability of the Helstrom bound we now consider a broader range of possible measurements. For discriminating multipartite quantum states, several studies have investigated the role of measurements beyond simply LOCC operations, namely separable and positive partial transpose (PPT) measurements [56, 57, 58, 59, 60]. Consider a $d$-dimensional quantum state shared between two parties, Alice and Bob. We shall denote the dimensions of the quantum state held by Alice and Bob as $d_{A}$ and $d_{B}$ respectively. Then a POVM $\{\Pi_{k}\}$ is separable if each $\Pi_{k}$ can be written as
$\Pi_{k}=\Pi_{k,A}\otimes\Pi_{k,B}\;,$ (25)
where $\Pi_{k,A}$ and $\Pi_{k,B}$ represent valid POVMs on Alice and Bob’s respective Hilbert spaces. It is clear that every LOCC measurement can be written as a separable measurement, however the converse is not necessarily true.
PPT measurements are a broader class of measurement again. The partial transpose operation acting on $A\otimes B$ takes the transpose of one of the subsystems, $A^{T}\otimes B$. PPT measurements are POVMs that remain positive under the action of partial transpose. The set of all PPT measurements is a closed convex cone, meaning that the optimal PPT measurement for state discrimination can be found via a semi-definite program (SDP) [61, 62]. Interestingly, SDPs appear in many other branches of quantum information [63, 64, 65, 66, 67].
Every separable measurement is a PPT measurement but the converse is not true. Therefore, we can construct a broader class of measurements where we split the $M$-copy quantum state to be discriminated into two groups of $M_{1}$ and $M_{2}$ copies, i.e. $d_{A}=d^{\otimes M_{1}}$ and $d_{B}=d^{\otimes M_{2}}$, where $d$ is the dimension of a single copy of the quantum state. We then allow PPT measurements between these two groups, and we denote the error probability for this strategy as $P_{M_{1},M_{2}}^{M,\text{ PPT}}$. Owing to the symmetry across the partition $P_{M_{1},M_{2}}^{M,\text{ PPT}}=P_{M_{2},M_{1}}^{M,\text{ PPT}}$. The results for this class of measurement are shown in Fig. 5. As in the previous section we see that even PPT measurements among any bipartition is not sufficient to saturate the $M$-copy Helstrom bound. However, it should be noted that, as expected, the PPT measurement performs better than the corresponding one-way LOCC measurement.
IV.5 Random examples to investigate the attainability of the Helstrom bound
Given any multi-qubit or multi-copy quantum state from a known set of quantum states to be discriminated we can efficiently compute the optimal PPT measurement given any partition. This is true even when discriminating more than two unknown quantum states. The Helstrom bound for discriminating many different states can also be written as an SDP [68]. This enables us to compare the PPT measurements to the optimal Helstrom measurement for many different random examples. This is done in Fig. 6 for two different settings. In the left figure we compare PPT measurements corresponding to the $P_{2,2}^{4,\text{ PPT}}$ and $P_{1,3}^{4,\text{ PPT}}$ strategies for discriminating between four copies of three known qubit states, i.e we distinguish between $\rho_{1}^{\otimes 4}$, $\rho_{2}^{\otimes 4}$ and $\rho_{3}^{\otimes 4}$. In this example, we find that in some situations the PPT measurement equals the Helstrom measurement to within the numerical accuracy of the SDP. In the right figure, we show the same, however now we consider discriminating between three unknown 16 dimensional quantum states. In this case, we do not find any examples where the PPT measurement matches the performance of the Helstrom measurement. However, this difference may be attributable to the fact that a 16 dimensional Hilbert space is a lot richer than the space formed by taking the tensor product of four qubit states. Note that for these random examples we do not assume that the states are equally likely, i.e. we do not assume a uniform prior distribution.